OpenStudy (anonymous):

Evaluate the following expression using the values given. Find x − 3y if x = −3 and y = −2

OpenStudy (callisto):

x - 3y Put (-3, -2) into the expression x-3y = (-3) - 3(-2) = -3 + 6 =?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow -3 - 3(-2) }\) \(\Large \color{MidnightBlue}{\Rightarrow -3 - (-6) }\) \(\Large \color{MidnightBlue}{\Rightarrow -3 + 6 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3 }\)

OpenStudy (anonymous):

thank you so much